# Integrated Communication, Navigation, and Remote Sensing in LEO Networks with Vehicular Applications

### v1

@@ Abstract
Traditionally, communication, navigation, and remote sensing (CNR) satellites are separately performed, leading to resource waste, information isolation, and independent optimization for each functionality. Taking future automated driving as an example, it faces great challenges in providing high-reliable and low-latency lane-level positioning, decimeter-level transportation observation, and huge traffic sensing information downloading. To this end, this article proposes an integrated CNR (ICNR) framework based on low earth orbit (LEO) satellite mega-constellations from the perspective of vehicular applications. After introducing the main working principles of the CNR functionalities to serve as the technological basis, we characterize the potentials of the integration gain in vehicular use cases. Then, we investigate the ICNR framework in different integration levels, which sheds strong light on qualitative performance improvement by sophisticatedly sharing orbit constellation, wireless resource, and data information towards meeting the requirements of vehicular applications. We also instantiate a fundamental numerical case study to demonstrate the integration gain and highlight the main tradeoffs in managing the ICNR networks from the perspective of vehicular applications.
@@ I Introduction
While deciding the altitude of a single satellite, there is a fundamental tradeoff between coverage range and signal propagation latency. However, by developing a huge constellation of low-altitude satellites, the coverage can be enhanced while reducing latency, which brings about a rising tide of world-wide LEO satellite network construction in the past ten years. Meanwhile, technological advancements have made it possible to build small cost-effective LEO satellites, develop reusable rockets, and adjust moving status in orbit. In this context, the economically feasible LEO satellite manufacture, launching, and in-orbit management techniques have enabled many countries, organizations, and commercial companies to establish their own satellite constellations. Besides the advantage of low latency, the huge constellation of LEO satellites is also good at enhancing capacity and reliability of communication, accuracy and rapidity of positioning, and spatial and temporal resolution of remote sensing [2]. Due to such a great potential, the number of LEO satellites may persistently grow towards the capacity of the orbits, predicted as several millions, leading to mega-constellations of LEO satellites in the future [3].
It had long been the situation that different kinds of satellites usually serve the customers with distinct purposes. Particularly, communication satellites mainly provide special users connectivity in regions out of terrestrial network coverage, such as sea, forest, and desert areas. Navigation satellites offer global positioning, navigation, and timing (PNT) services to a large body of location-dependent applications, such as object tracking, route planning, and industrial automation. Remote sensing satellites are able to provide a top view of the earth, which can be applied to agriculture monitoring, natural disaster discovering, and traffic network optimization. However, heterogeneous services are required by vehicular terminals, especially in the era of autonomous driving. For example, an automated connected vehicle in intelligent transportation systems may require timely remote-sensing traffic information to optimize its route planning, necessitate high-precision positioning to perform lane-level navigation, and need high-capacity communication to transfer diversified messages including remote sensing data [4]. To provide the required CNR services to future vehicular applications, the LEO mega-constellation of satellites is expected to be a promising solution.
@@
Traditional separate-deployed CNR satellite systems face considerable challenges in meeting captious vehicular requirements due to loss of sharing gain in terms of orbit constellation, wireless resource, and data information, which further prohibits joint CNR optimization towards satisfying the autonomous driving application. In this regard, the framework of integrated communication and navigation has been proposed. In the early stage of such integration dating back to 1980s, high-altitude satellites are mainly designed for vehicular navigation, implying that navigation is taken as the primary function and communication is introduced as the secondary capability. With the proliferation of LEO satellite based broadband Internet, the integration moves the foundation to the communication side, where navigation data is incorporated into the communication packets [5]. Another topic is the integration of communication and remote sensing, which has hardly been considered until the recent flourishing low-altitude space networks, mainly because the delay requirement of remote sensing data downloading for traditional users is not as strict as present applications, such as intelligent transportation. Particularly, communication among remote sensing satellites offers the opportunity of onboard sensing information fusion, leading to higher resolution and quicker response, which is pretty beneficial to modern vehicular users [6]. Other than the combination of dual functionalities, a few works have begun to consider the combination of the three functionalities. Frequency diverse array and orthogonal frequency division multiplexing have been utilized in [7] to efficiently exploit waveform diversity for integrated communication, sensing, and navigation. In [8], the constellation design has been investigated and a simplified analysis has been provided to jointly consider the communication and remote sensing requirements. The attractive prospect of the integration with applications in ecological environment monitoring has been demonstrated in [9]. Last but not least, the national aeronautics and space administration (NASA) has also raised its interests on multi-purpose space cubes with remote sensing, communication, and navigation integration [10].
@@
Despite sporadically mentioned in the literature, the integration of the CNR in a mega-constellation of LEO satellite networks has not been completely investigated with its fundamentals, classification, potentials, and challenges being specified. In this article, we first summarize the most fundamental principles of the CNR techniques individually, followed by demonstration of the proposed integrated CNR (ICNR) framework with vehicular applications as a representative use case. Then, different levels of integration in the proposed ICNR network are discussed and a fundamental numerical case study is conducted to demonstrate the integration gain. Finally, we highlight several main tradeoffs that need to be elaborated in practical ICNR network deployment and management.
@@ II-A Satellite Communication
Traditional communication satellites usually act as relays to transit information, i.e., forwarding information from a source node to a destination node on the earth, where the destination node can be as far as tens of thousands of kilometers away from the source node. Taking downlink transmission as an example, we consider terminal 0 as the reference user receiving information signal from satellite 0. Without loss of generality, there are additional $M$ satellites in the zenith of the terminal, causing interference to the terminal. Denote $x_{m}(t)$ as the transmitted signal from satellite $m$ at time $t$. Then, the received signal at time $t$ can be expressed as
$y_{0}(t)=h_{0}x_{0}(t-\tau_{0})+\sum_{m=1}^{M}h_{m}x_{m}(t-\tau_{m})+n(t),$ (1)
where $n(t)$ is the receiving noise, and $\tau_{m}$ and $h_{m}$ are the propagation delay and channel coefficient from satellite $m$ to the terminal, respectively. The main task of satellite communication is to recover $x_{0}(t)$ from $y_{0}(t)$ by removing the interference and noise, i.e., the second and third terms in (1), respectively. Generally, capacity, reliability, and latency are the three dominant metrics that are usually cared about by vehicular users, e.g., to download the remote sensing data. A peculiarity of vehicular applications is the mobility of terminals, which makes channel features to be different from static users and thus deserves special attention in practical systems.
@@ II-B Satellite Navigation
Positioning and timing are the two main tasks of navigation systems, which are the most fundamental requirements of vehicular terminals. Since the ephemeris and almanac of satellites are available and atomic clocks are equipped in satellites, navigation satellites can be aware of their positions as well as time. Consider that a target user aims to determine the time $t$ and its position denoted by three-dimensional coordinates $(u,v,w)$ in a Cartesian coordinate system. To this end, upon the user receiving a signal from satellite $k$ emitted at time $t_{k}$, it is able to get the distance from itself to satellite $k$ being $c(t-t_{k})$, where $c$ is the speed of light and $t_{k}$ is the time at which the signal is emitted by satellite $k$. Using simple trigonometry, we have the equation given by
$(u-u_{k})^{2}+(v-v_{k})^{2}+(w-w_{k})^{2}=c^{2}(t-t_{k})^{2},$ (2)
where $(u_{k},v_{k},w_{k})$ is the position of satellite $k$. Since the user needs to determine four unknown parameters, $u$, $v$, $w$, and $t$, at least four satellites are required to establish four equations as in (2) for $k=1,2,3,4$. Due to ranging errors caused by inaccurate reference satellite position, multi-path propagation, light refraction, etc., the position and time are estimated rather than computed exactly. Generally, receiving signals from additional satellites does positioning and timing a favor with more precision by finding a least-square solution of (2) with more than four available navigation satellites. Besides the number of visible satellites, the relative geometric relationship among user equipment and visible satellites also plays a significant role in determining the accuracy of positioning and timing, usually measured by the geometric dilution of precision (GDOP) indicating the ratio of positioning and timing error to ranging error [11]. Vehicular terminals can also leverage inertial measurement units (IMU) and cooperative technique to improve the positioning accuracy in addition to using the available satellite navigation signals.
@@ II-C Satellite Remote Sensing
Remote sensing satellites are designed to view the earth from a position in orbit and usually produce images of the earth surface. These images are crucial to a diverse array of applications, including environmental monitoring, weather prediction, agriculture management, disaster detection, and intelligent transportation. There are generally two kinds of satellite remote sensing technologies, namely passive and active ones. The passive remote sensing satellites detect energy emitted or reflected from objects on the earth, usually operating in visible, infrared, thermal infrared, and microwave portions of the electromagnetic spectrum. Since the passive sensing technology is vulnerable to lighting conditions and dependent on ambient energy, the active remote sensing, especially synthetic aperture radar (SAR) based remote sensing, plays an important role in mission-critical applications. Similar to general radar remote sensing mechanism, a SAR satellite performs electromagnetic radiations to illuminate the target on the earth and subsequently records the reflected radiation signal to get earth observations. This implies that it is difficult for a remote sensing satellite to achieve high-resolution images by very narrow beam focus due to antenna size and working frequency limitations. However, by exploiting the motion of spacecrafts, SAR satellites can achieve equivalent long-aperture antennas by accumulating signal reception over time, leading to an azimuth resolution $\delta_{\rm a}=\frac{d_{\rm a}}{2}$, where $d_{a}$ is the aperture length of the antenna in the azimuth direction. Although a SAR satellite prefers a small $d_{a}$, practical antenna size should be large enough to restrict the illuminated ground area to avoid ambiguous echo signal reception [12]. With the pulse compression technique, the cross-track resolution of a SAR satellite can be estimated as
$\delta_{\rm c}=\frac{c}{2B\sin\varphi},$ (3)
where $B$ is the spectrum bandwidth and $\varphi$ is the view angle, i.e., the satellite centered angle between the ground target and the nadir. In vehicular use cases, the spatial resolution should be low enough to distinguish different vehicles. Besides the spatial precision, the temporal precision also deserves special attention, which is traditionally captured by the revisit time. It is worth noting that the temporal resolution should be defined with respect to the whole system in a mega-constellation of LEO satellites, i.e., the time it takes for a target to be observed twice by the constellation rather than by a satellite. The temporal resolution needs to be as small as the variation period of the road traffic in vehicular applications, which can be measured by age of information (AoI) of the remote sensing data received by vehicles [13].
Figure 3: Illustration of the different integration levels of the proposed ICNR framework.
@@ II-D Proposed ICNR Framework
Particularly, operating in a mega-constellation LEO satellite networks, the ICNR framework sheds new light on rural and urban vehicular applications, as illustrated in Fig. 2. In case of being trapped on roads due to natural diaster in distant area without terrestrial infrastructure coverage, the vehicle can first get its enhanced localization by exploiting more visible LEO satellites and then send an save-our-ship (SOS) signal to an available satellite, which further forwards the SOS signal to a ground rescue station. On the other hand, an appropriate satellite can be triggered to perform real-time remote sensing for the related region and then send the disaster sensing information to the ground rescue station. Based on the geographical disaster condition and the location of the stranded vehicle, the best rescue plan can be made, e.g., determining the best transportation means and the rescue trajectories.
@@
The ICNR networks are also quite beneficial to future intelligent transportation systems (ITS) even within the coverage of terrestrial networks, as presented in the urban area in Fig. 2. Because of stochastic traffic distribution, the shortest driving path may not be the best choice for a vehicle. In this context, a satellite may first implement remote traffic sensing for the area covering both the current location and the destination point. Then, such sensed traffic information is forwarded through ISLs to other satellites, which are closer to the vehicle or the ITS scheduling center. After that, a proper relaying satellite can transfer the traffic distribution information to the vehicle to let it make a better route plan. Alternatively, a relaying satellite may also send the traffic distribution information to the ITS scheduling center, which produces a route suggestion after considering the whole traffic system and notifies the vehicle. On the other hand, it becomes feasible to achieve more enhanced centimeter-level positioning by exploiting more visible LEO satellites and simultaneous multi-purpose signal transmission. Finally, based on the timely traffic congestion information, the accurate positioning, and the suggestion from the ITS center, the vehicle is able to make the optimal route decision and lane-level navigation. In summary, real-time vehicle-distinguishable remote sensing, lane-level positioning, and huge remote sensing information downloading are commonly required by intelligent connected vehicles, which may be suitably addressed by the proposed ICNR networks with a mega-constellation of LEO satellites.
@@ III Integration Levels of ICNR
Depending on the coupling degrees among different functionalities, we classify the proposed ICNR framework into four stages, i.e., function-level, signal-level, data-level, and application-level integration, in an increasing order of the coupling degree, as summarized in Fig. 3.
@@
Function integration stands for the lowest integration level of the ICNR networks, where the CNR functionalities are loaded in a single LEO satellite but operate almost independently. Each functionality uses its exclusive hardware and software components to accomplish its own task, except some underlying equipment, such as the battery system and central on-board processing unit that are shared by the CNR. This means that the ICNR networks can be sometimes treated as three mutually independent systems for the CNR, with a common constellation shared. As the equivalent number of satellites for each functionality increases by such integration, the CNR performance can be overall improved in terms of communication coverage, positioning resolution, and remote sensing response time, etc. A typical potential of such integration is to leverage the broadband communication capability to perform resolution enhanced multi-satellite sensing fusion and low-latency transportation sensing information downloading for automated driving applications.
@@
In this level, the signals for the CNR are expected to be combined, i.e., the signal emitted by the satellites carries communication information, navigation reference, and remote sensing probes simultaneously, which leads to a more efficient use of limited spectrum resource than the function integration. We can also have an easier variant that flexibly reuses the wireless resource in time, frequency, or space domains. Generally, there is a fundamental tradeoff among different performance of the CNR in designing multi-purpose signals. In practical systems, how functionalities are combined in the signal should depend on the realtime geographical distribution of the CNR requirements. Particularly, for a time instance at which only communication and navigation services are required by the ground users in a region, there is no need to embrace the remote sensing function in the current signal design. The vehicular terminals can benefit greatly from equivalent enlarged resource pool with signal-level integration, which holds attractive potentials to provide high-resolution traffic sensing information and lane-level navigation.
@@
Different from the above two levels, data-level integration allows the satellites to share information collected by different functionalities, where the information coming originally from one functionality may be very helpful in optimizing the network performance from the perspective of another functionality. The remote sensing results generally show the geographical vehicular traffic distribution, which can be leveraged to conduct more efficient traffic-aware communication and navigation resource scheduling. Using navigation information, such as location and speed of vehicles, the communication channel may be more efficiently predicted, which allows more adaptive communication transceiver design. By analyzing the communication channel state information (CSI) of a large number of vehicular users, we can estimate the overall channel conditions for a given region, which can be exploited to perform better remote sensing beamforming design and CSI fingerprinting based positioning. Finally, high mobility of ground vehicles causes rapid change of remote transportation sensing images, navigation GDOP values, and communication load distribution, which gives rise to a high demand on fast inter-functionality data mining.
@@
All information delivery, localization, and earth surface observation functionalities are actually supporting technologies rather than the final objectives of the ICNR system. To this end, the application-level integration aims to achieve the ultimate goal on providing satisfied services for various applications of ground users. As previously mentioned, an autonomous driving application simultaneously requires the CNR services. Therefore, the application-level integration takes the waveform design, resource allocation, and data exploitation as a whole to jointly optimize all procedures, as illustrated in Fig. 3. Particularly, coordination and synchronization among different functionalities within each satellite and even among different satellites are particularly needed to avoid interference and conflict. In such cases, providing on-demand services is one of the key features of the ICNR networks, implying that an increasing demand on one of the three functionalities triggers the system resource allocation tilting to it. As future vehicular applications often require comprehensive CNR services for autonomous driving, high reliability and low latency are usually required in communication, positioning, and remote transportation sensing. The goal-oriented application-level integration holds great potentials to jointly satisfy such heterogeneous requirements by targeted resource scheduling.
@@
Requirement/Parameter Value or Description
[$M_{c},M_{n},M_{r}$] $[4500,250,250]$
[$B_{c},B_{n},B_{r}$] $[250,25,25]$ MHz
@@
Large-scale fading $110+37.6\log_{10}(d)$

### v2

@@ Abstract
Traditionally, communication, navigation, and remote sensing (CNR) satellites are separately performed, leading to resource waste, information isolation, and independent optimization for each functionality. Taking future automated driving as an example, it faces great challenges in providing high-reliable and low-latency lane-level positioning, decimeter-level transportation observation, and huge traffic sensing information downloading. To this end, this article proposes an integrated CNR (ICNR) framework based on low Earth orbit (LEO) satellite mega-constellations. After introducing the main working principles of the CNR functionalities to serve as the technological basis, we characterize the potentials of the integration gain in vehicular use cases. Then, we investigate the ICNR framework in different integration levels, which sheds strong light on qualitative performance improvement by sophisticatedly sharing orbit constellation, wireless resource, and data information towards meeting the requirements of vehicular applications. We also instantiate a fundamental numerical case study to demonstrate the integration gain and highlight possible future research directions in managing the ICNR networks.
@@ I Introduction
While deciding the altitude of a single satellite, there is a fundamental tradeoff between coverage area and signal propagation latency. However, by developing a huge constellation of low-altitude satellites, the coverage can be enhanced while reducing latency, which brings about a rising tide of world-wide LEO satellite network construction in the past ten years. Meanwhile, technological advancements have made it possible to build small cost-effective LEO satellites, develop reusable rockets, and adjust moving status in orbit. In this context, the economically feasible LEO satellite manufacture, launching, and in-orbit management techniques have enabled many countries, organizations, and commercial companies to establish their own satellite constellations. Besides the advantage of low transmission latency and small signal attenuation, the huge constellation of LEO satellites is also good at enhancing capacity and reliability of communication, accuracy and rapidity of positioning, and spatial and temporal resolution of remote sensing [2]. Due to such a great potential, the number of LEO satellites may persistently grow towards the capacity of the orbits, predicted as several millions, leading to mega-constellations of LEO satellites in the future [3].
It had long been the situation that different kinds of satellites usually serve the customers with distinct purposes. Particularly, communication satellites mainly provide special users connectivity in regions out of terrestrial network coverage, such as sea, forest, and desert regions. Navigation satellites offer global positioning, navigation, and timing (PNT) services for a large body of location-dependent applications, including outdoor expedition, drone management, and industrial automation. Remote sensing satellites are able to provide a top view of the earth, which can be applied to agriculture monitoring, natural disaster discovering, and traffic network optimization. However, heterogeneous services can be required by vehicular terminals, especially in the era of autonomous driving. In particular, an automated connected vehicle in intelligent transportation systems may require timely remote-sensing traffic information to optimize its route, necessitate high-precision positioning to perform lane-level navigation, and need high-capacity communication to transfer diversified messages including remote sensing data [4]. To efficiently provide comprehensive CNR services, the LEO mega-constellation of satellites is expected to be a promising solution.
@@
Traditional separately deployed CNR satellite systems face considerable challenges in meeting captious vehicular requirements due to loss of sharing gain in terms of orbit constellation, wireless resource, and data information, which further prohibits joint CNR optimization towards satisfying the autonomous driving application. In this regard, the framework of integrated communication and navigation has been proposed. In the early stage of such integration dating back to 1980s, high-altitude satellites are mainly designed for vehicular navigation, implying that navigation is taken as the primary function and communication is introduced as the secondary capability. With the proliferation of LEO satellite based broadband Internet, the integration moves the foundation to the communication side, where navigation data is incorporated into the communication packets [5]. Another topic is the integration of communication and remote sensing, which has hardly been considered until the recent flourishing low-altitude space networks, mainly because the delay requirement of remote sensing data downloading for traditional users is not as strict as present applications, such as intelligent transportation. Particularly, communication among remote sensing satellites offers the opportunity of onboard sensing information fusion, leading to higher resolution and quicker response, which is pretty beneficial to modern vehicular users [6]. Other than the combination of dual functionalities, a few works have begun to consider the combination of the three functionalities. Frequency diverse array and orthogonal frequency division multiplexing have been utilized in [7] to efficiently exploit waveform diversity for integrated communication, sensing, and navigation. In [8], the constellation design has been investigated and a simplified analysis has been provided to jointly consider the communication and remote sensing requirements. The attractive prospect of the integration with applications in ecological environment monitoring has been demonstrated in [9]. Last but not least, the national aeronautics and space administration (NASA) has also raised its interests on multi-purpose space cubes with remote sensing, communication, and navigation integration [10].
@@
Despite sporadically mentioned in the literature, the integration of the CNR in a mega-constellation of LEO satellite networks has not been completely investigated with its fundamentals, classification, potentials, and challenges being specified. In this article, we first summarize the most fundamental principles of the CNR techniques individually, followed by demonstration of the proposed integrated CNR (ICNR) framework with vehicular applications as a representative use case. Then, different levels of integration in the proposed ICNR network are discussed and a fundamental numerical case study is conducted to demonstrate the integration gain. Finally, we touch on possible future research directions that need to be elaborated in practical network deployment and management.
@@ II-A Satellite Communication
Traditional communication satellites usually act as relays to transit information, i.e., forwarding information from a source node to a destination node on the earth, where the destination node can be as far as thousands of kilometers away from the source node. Taking downlink transmission as an example, we consider a terminal trying to capture information signal, $x_{0}(t)$, from a host satellite in Fig. 1. In the zenith of the terminal, there usually exist some additional satellites emitting specified signals, such as $x_{1}(t)$ and $x_{2}(t)$, to their associated users, causing interference to the terminal. In this regard, the terminal generally receives delayed and attenuated signals from the host satellite and the interfering satellites as well. The main task of satellite communication is to recover the target signal from the superimposed signal by removing the interference and noise. Generally, capacity, reliability, and latency are the three dominant metrics that are usually cared about by vehicular users, e.g., to download the remote sensing data. A peculiarity of vehicular applications here is the mobility of terminals, which makes channel features to be different from static users and deserves special attention in practical systems.
@@ II-B Satellite Navigation
Positioning and timing are the two main tasks of navigation systems, which are the most fundamental requirements of vehicular terminals. Navigation satellites can usually be aware of time and their trajectories, with the aid of clocks equipped and periodical ephemeris acquisition from ground observation stations. Consider that a target user aims to determine the time $t$ and its position denoted by three-dimensional coordinates $(u,v,w)$ in a Cartesian coordinate system. To this end, upon the user receiving a signal from a reference satellite, it is able to get the distance from itself to the reference satellite by multiplying the speed of light and the time it takes for the signal to travel. Such a distance can also be expressed by calculating the Euclidean distance between the positions of the target user and the reference satellite using simple geometry. Then, with one reference satellite in hand, we have one distance equation with four unknown variables. To determine these four unknowns, at least four reference satellites are required to establish four independent distance equations, given that the position and clock time of the $i$th reference satellite, jointly denoted by $(u_{i},v_{i},w_{i},t_{i})$, are konwn. Due to ranging errors caused by inaccurate reference satellite position, multi-path propagation, light refraction, etc., the position and time are estimated rather than computed exactly. Generally, receiving signals from additional satellites does positioning and timing a favor with more precision by finding a least-square solution of the distance equations with more than four available navigation satellites. Besides the number of visible satellites, the relative geometric relationship among user equipment and visible satellites also plays a significant role in determining the accuracy of positioning and timing, usually measured by the geometric dilution of precision (GDOP) indicating the ratio of positioning and timing error to ranging error [11]. Vehicular terminals can also leverage inertial measurement units (IMU) and cooperative technique to improve the positioning accuracy in addition to exploiting available satellite navigation signals.
@@ II-C Satellite Remote Sensing
Remote sensing satellites are designed to view the earth from a position in orbit and usually produce images of the earth. These images are crucial to a diverse array of applications, including environmental monitoring, weather prediction, agriculture management, disaster detection, and intelligent transportation. There are generally two kinds of satellite remote sensing technologies, namely passive and active ones. The passive remote sensing satellites detect energy emitted or reflected from objects on the earth, usually operating in visible, infrared, thermal infrared, and microwave portions of the electromagnetic spectrum. Since the passive sensing technology is vulnerable to lighting conditions and dependent on ambient energy, the active remote sensing, especially synthetic aperture radar (SAR) based remote sensing, plays an important role in mission-critical applications. Similar to general radar remote sensing mechanism, a SAR satellite performs electromagnetic radiations to illuminate the target on the earth and subsequently records the reflected radiation signal to get earth observations. Due to limitations on antenna size and waveform frequency, it is generally difficult to capture high-resolution images of the surface of Earth. However, by exploiting the motion of spacecrafts, SAR satellites can achieve equivalent long-aperture antennas by accumulating signal reception over time, leading to a resolution to be a half of the antenna length in the azimuth direction. Although a SAR satellite prefers a small length of the antenna, practical antenna size should be large enough to restrict the illuminated ground area to avoid ambiguous echo signal reception [12]. With the pulse compression technique, the cross-track resolution of a SAR satellite is proportional to the speed of light divided by the product of the spectrum bandwidth and the sine of the view angle, where the view angle is the satellite centered angle between the ground target and the nadir. In vehicular use cases, the spatial resolution should be high enough to distinguish different vehicles. Besides the spatial precision, the temporal precision also deserves special attention, which is traditionally captured by the revisit time. It is worth noting that the temporal resolution should be defined with respect to the whole system in a mega-constellation of LEO satellites, i.e., the time it takes for a target to be observed twice by the constellation rather than by a satellite. The temporal resolution needs to be comparable to the variation period of the road traffic in vehicular applications, which can be measured by age of information (AoI) of the remote sensing data received by vehicles [13].
Figure 3: Key components and working mechanisms of four different integration levels of the proposed ICNR framework.
@@ II-D Proposed ICNR Framework
Particularly, operating in a mega-constellation LEO satellite networks, the ICNR framework sheds new light on rural and urban vehicular applications, as illustrated in Fig. 2. In case of being trapped on roads due to natural diaster in distant regions without terrestrial infrastructure coverage, the vehicle can first get its enhanced localization by exploiting more visible LEO satellites and then send a save-our-ship (SOS) signal to an available satellite, which further forwards the SOS signal to a ground rescue station. On the other hand, an appropriate satellite can be triggered to perform real-time remote sensing for the related region and then send the disaster sensing information to the ground rescue station. Based on the geographical disaster condition and the location of the stranded vehicle, the best rescue plan can be made, e.g., determining the best transportation means and the rescue trajectories. The ICNR system is also expected to perform remote driving operation monitoring and cooperative perception in rural areas with the aid of high-bandwidth satellite communications.
@@
The ICNR networks are also quite beneficial to future intelligent transportation systems (ITS) even within the coverage of terrestrial networks, as presented in the urban area in Fig. 2. Because of stochastic traffic distribution, the shortest driving path may not be the best choice for a vehicle. In this context, a satellite may first implement remote traffic sensing for the area covering both the current location and the destination point. Then, such sensed traffic information is forwarded through ISLs to other satellites, which are closer to the vehicle or the ITS scheduling center. After that, a proper relaying satellite can transfer the traffic distribution information to the vehicle to let it make a better route plan. Alternatively, a relaying satellite may also send the traffic distribution information to the ITS scheduling center, which produces a route suggestion after considering the whole traffic system and notifies the vehicle. Although terrestrial ITSs can deploy a massive number of sensors on roads, it is quite difficult for vehicles to periodically get synchronous traffic information from these sensors due to limited communication resource. Even such information can be timely collected, it is still challenging to depict a whole picture of the road traffic because heterogeneous sensors produce multi-mode data and different equipment have different angles of view. By contrast, the proposed ICNR system is able to observe the interested transportation area as a whole and outputs a high-resolution remote sensing image, which can be downloaded to vehicles by high-capacity communication. Furthermore, aircrafts in the low sky, e.g., unmanned aerial vehicles (UAVs), can also be deployed as complementary platforms to perform remote traffic sensing for the regions that are not directly visible to satellites. Other than the traffic distribution information acquisition, it becomes feasible for the ICNR system to achieve more enhanced centimeter-level positioning by exploiting more visible LEO satellites and simultaneous multi-purpose signal transmission. Finally, based on the timely traffic congestion information, the accurate positioning, and the suggestion from the ITS center, vehicles are able to make the optimal route decision and lane-level navigation. In summary, real-time vehicle-distinguishable remote sensing, lane-level positioning, and huge remote sensing information downloading are commonly required by intelligent connected vehicles, which may be suitably addressed by the proposed ICNR networks with a mega-constellation of LEO satellites. Particularly, the superiority becomes more pronounced in the case of ground network congestion and terrestrial emergency even in urban areas.
@@ III Integration Levels of ICNR
In this section, we classify the proposed ICNR framework into four stages, i.e., function-level, signal-level, data-level, and application-level integration, in an increasing order of the coupling degree, where key system components and working mechanism of different integration levels are summarized in Fig. 3.
@@
Function integration stands for the lowest integration level of the ICNR networks, where the CNR functionalities are loaded in a single LEO satellite but operate almost independently. Each functionality uses its exclusive hardware to accomplish its own task, except sharing the same orientation and attitude with respect to each satellite. This means that the ICNR networks can be sometimes treated as three mutually independent systems for the CNR, with a common constellation shared. The key system module in such a level of integration is the platform management, including components of hardware requirement analysis, inter-functionality coordination, and finally in-orbit satellite control. As the equivalent number of satellites for each functionality increases by such integration, the CNR performance can be overall improved in terms of communication coverage, positioning accuracy, and remote sensing response time. A typical potential of such integration is to leverage the broadband communication capability to perform resolution enhanced multi-satellite sensing fusion and low-latency transportation sensing information downloading for automated driving applications.
@@
In this level, the signals for the CNR are expected to be combined, i.e., the signal emitted by a satellite carries communication information, navigation reference, and remote sensing probes simultaneously, which leads to a more efficient use of limited spectrum resource than the function-level integration. As a core system module, waveform fusion can be achieved by channel environment estimation and multipurpose beam design, based on the platform management as demonstrated in the function-level integration. We can also have an easier variant that allocates different functionalities orthogonal wireless resource by division of time, frequency, or space. Generally, there is a fundamental tradeoff among different performance of the CNR in designing multi-purpose signals. In practical systems, how functionalities are combined in the signal should depend on the real-time geographical distribution of the CNR requirements. Particularly, for a time instance at which only communication and navigation services are required by the ground users in a region, there is no need to embrace remote sensing capability in the current signal design. The vehicular terminals can benefit greatly from equivalent enlarged resource pool with signal-level integration, which holds attractive potentials to provide high-resolution traffic sensing information and lane-level navigation.
@@
Different from the above two levels, data-level integration allows the satellites to share information collected by different functionalities, where the information coming originally from one functionality may be very helpful in optimizing the network performance from the perspective of another functionality. The most prominent system module here is the information sharing, which is composed of three leading components, i.e., CNR data mining, knowledge based optimization, and waveform fusion. For instance, the remote sensing results generally show the geographical vehicular traffic distribution, which can be leveraged to conduct more efficient traffic-aware communication and navigation resource scheduling. Using navigation information, such as location and speed of vehicles, the communication channel may be more efficiently predicted, which allows more adaptive communication transceiver design. By analyzing the communication channel state information (CSI) of a large number of vehicular users, we can estimate the overall channel conditions for a given region, which can be exploited to perform better remote sensing beamforming design and CSI fingerprinting based positioning. Finally, high mobility of ground vehicles causes rapid change of remote transportation sensing images, navigation GDOP values, and communication load distribution, which gives rise to a high demand on fast inter-functionality data mining.
@@
All information delivery, terminal localization, and earth observation functionalities are actually supporting technologies rather than the final objectives of the ICNR system. To this end, the application-level integration aims to support vehicular tasks, rather than meet predefined CNR requirements as in the other three levels of integration. In particular, the CNR service requirements raised by a specific application can be expressed as a multi-dimensional domain, where different performance aspects can be somewhat interchangeable. Taking autonomous driving as an example, the vehicle may require periodic high-resolution traffic sensing data to perform adaptive route planning, or alternatively more frequent acquisition of traffic information with less accuracy, i.e., the requirements on communication and remote sensing can be interchangeable. Based on the real-time system condition, such as satellite geometry, channel environment, and user distribution, the application-level integration can choose an appropriate service target within the CNR demand domain, which will be provided by intelligent global optimization with cross-layer joint consideration of information sharing, waveform fusion, and platform management. Therefore, towards supporting various vehicular applications, the system requires a task-oriented service providing module, consisting of service demand domain analysis, network condition evaluation, and intelligent global optimization components. The intelligent service supply adaptive to demand domain and network condition holds great potentials to satisfy a massive number of vehicular users with diversified applications.
@@
Requirement/Parameter/Scenario Value or Description
Numbers of traditional CNR satellites $4500$, $250$, $250$
Bandwidth of traditional CNR satellites $250$, $25$, $25$; MHz
@@
Driving area Remote rural region without terrestrial coverage
Vehicle density Very sparse without inter-vehicle connection
Velocity of vehicles Less than 108 km/h
Perception and maneuver reporting rate Once every 3 meters
Perception data size 100 detected objects with 80 bytes for each [15]
Maneuver data size 500 bytes [15]
@@
Large-scale fading $110+37.6\log_{10}({\rm distance})$, ${\rm distance}$ in km